# Scaling-aware rating of count forecasts

### v4

@@ 5 Application to Demand Edge for Retail predictions at Sainsbury’s
The M5 dataset discussion in the previous section has the great advantage to be reproducible by anyone. On the other hand, one might argue that the setup is artificial, since one deals with a curated dataset that does not contain all data that are used in production (e.g. the price of the product). To complement our discussion and prove the applicability of our method in practice, we apply the scaling-aware rating to the predictions of the Demand Edge for Retail forecasts produced by Blue Yonder for the UK retailer Sainsbury’s. Demand Edge for Retail is based on a causal machine-learning model that computes the contribution of features such as price, weather, day-of-the-week, holiday, promotion and many others to the expected demand of a product (Wick et al., 2019, 2021). It produces a forecast that consists of a prediction for the expected mean as well as for the variance, technical details can be found in (Wick, a, b, c). That is, the prediction itself estimates the super-Poissonian uncertainty. To simplify the discussion and make it comparable to the M5 dataset discussion, we only use the predicted mean, and interpret it as predicted mean of the Poisson distribution.
@@
We analyse the next-day-forecasts (horizon 1) for Monday, October 2nd to Saturday, October 28th, 2023, excluding Sundays. We exclude items that are not sold in units of pieces but volume (liters) or mass (KGs), predictions that are smaller than 0.1 / product-location-day, and unsold product-day-location combinations for which the prediction was 20 or larger. The latter filter is motivated as follows: A Poisson prediction of $\mu=20$ forecasts the probability to sell 0 items to be about $2\cdot 10^{-9}$. Thus, even if the model were grossly overestimating, a sales value of 0 is absolutely unexpected and clearly indicates an out-of-stock situation. After data cleaning, we remain with a total of more than 250 million product-location-day combinations, corresponding to more than 500 million sold items. The global over-prediction of the model is less than 1%.
@@
Figure 10 shows the bucketed scatter plots for bias and noise for the Sainsbury’s dataset. It confirms the calibration of the forecast (outside the ultra-slow-movers), and the only slightly super-Poissonian noise over three orders of magnitude.
The direct comparison of scaling-unaware metrics would not be meaningful Kolassa (2008), but we can compare Figures 7 and 10 to perform a hybrid comparison: We see how the Demand Edge for Retail model outperforms the M5 competition models in the sense that, for a given rate bucket, it produces lower values in NMRPS. This is also reflected by the excellent values $S_{\text{overall}}\text{NMRPS}=84.7\%$ and $S_{\text{overall}}\text{Bias}=78.7\%$ on the Sainsbury’s dataset.
@@
Our approach hence confirms that, on a global level, the prediction is comparable to what could possibly be expected from first principles (Tichy, 2023c), as discussed in Section 2.1. A more in-depth investigation confirms that the reasons for non-Poissonian behavior described in Section 2.2 above are indeed relevant in this scenario (there are, for example, substantial “multi-buy” events when customers buy several pieces of one item). A thorough quantitative understanding of the mild departure from the Poisson ideal case based on those effects would be highly appreciated, but remains beyond the scope of the current work.
@@
As a more in-depth analysis, one can segregate data by different dimensions and ask questions as the following: Are certain stores predicted better than others? Are certain product groups predicted better than others? Are certain days or weekdays predicted better than others? All of these comparisons risk to induce the naïve scaling-trap, which is resolved by our approach. As an instructive example, we can differentiate by the highest-level product group and focus on the three largest assortments (Fresh, Ambient, Non foods). The resulting metrics are shown in Table 2.
@@
Similarly to our artificial example earlier in Section 2.6, the naïve scaling trap is lurking. Large differences between the three product groups could lead an analyst to conclude a dramatically different forecasting performance. More insight can be gained via the prediction-bucketed representation in Figure 11: The upper panel reflects the low bias per prediction bucket (although predicted slow movers are slightly underestimated). The lower panel confirms that the differences in forecast quality between the product groups is small, the difference of achieved KPIs in Table 2 is mostly borne by the different selling velocities (Fresh sells at about twice the mean velocity, Non foods sells at about one third of the mean velocity). When asking for improvement potential, it is worth noting that Fresh is performing slightly better than Ambient and Non foods in the slow- to mid-mover regime (mean prediction between 1/day to 10/day). The outliers in Ambient for fast movers (mean prediction $>$100/day) are worth investigating, but constitute only a tiny fraction of Ambient product-location-days and sales. Since the circle size scales with the square-root of the number of measurements, small buckets are visually over-emphasized.
The bucketed representation is summarized by the following table, which contains the ratings regarding NMRPS and bias, using the same parameters as for the M5 competition to make the numbers comparable (Blue Yonder uses internally different sets of industry-specific parameters to define the quality grades):
@@
Figure 11: Prediction-bucketed representation of bias (upper panel) and noise (NMRPS, lower panel), for Demand Edge for Retail predictions by Blue Yonder for UK retailer Sainsbury’s, differentiated by the highest-level product group. Circle sizes reflect the square-root of total number of sales contained in the respective buckets.
@@
In conclusion, applying scaling-aware rating on different data-subsets such as departments, locations, days-of-the week helps focus development efforts on those specific forecasts that exhibit the largest gap to the Poisson ideal case. These are often not those that exhibit extremal values of scaling-unaware metrics! Scaling-aware rating accounts for the most important influence on metric values, the precision scaling of the Poisson distribution. It thereby avoids the naïve scaling trap so the user can focus on actual quality differences between models and between datasets.
@@ 6 Conclusions and outlook
Applied forecast judgement methodology should reconcile prevailing business practice (Carvallo, 2021; Bower, 2023) with the state of the art in probabilistic forecasting (Czado et al., 2009; Gneiting, 2011; Kolassa, 2016; Wei and Held, 2014; Gneiting and Katzfuss, 2014). Scaling-aware forecast rating provides a framework to tackle the problem of benchmarking and precision scaling in a systematic way. The counting statistics – the scaling of the Poisson distribution and of the noise of forecast quality references – can be modelled universally. This allows to separate these general effects and natural bounds from industry-specific aspects, which are captured by the parameters that define “excellent” to “unacceptable”. By adapting the parametrization introduced in Section 3.3, one can apply scaling-aware forecast rating to different forecasting problems, e.g. to industries in which bad data quality is common or certain influencing factors are known to be uncontrollable.
@@ 6.1 Applicability
Given a forecast, our method permits users to judge the status quo and whether desired improvements in accuracy are realistic at all: If forecast performance is close to the ideal Poisson prediction (or another distribution known to be the sharpest possible for the task at hand), the ask for a substantial improvement needs to be refused. If a gap to the Poissonian case emerges, the reason for this gap needs to be understood. Forecast creators should then elaborate whether it is an unrealistic expectation – the assumed “ideal” case is unreachable due to reasons of forecastability, or whether the forecast can actually be improved.
When several models compete, using scaling-aware methods makes their comparison more transparent and insightful: A given improvement of a metric can be understood holistically by analysing the respective bucket scatter plot as in Figures 7, 8, 10, 11. Two competing models can be placed onto a meaningful scale from “perfect” to “unacceptable”, providing context and orientation whether a certain improvement is substantial and solid, or little more than a mere coincidence.
The greatest added value of our method lies in dataset comparison: A comparison of the values of scaling-unaware metrics such as MAE, WMAPE or (N)MRPS for two sub-datasets of the same model is useless, since the naïve scaling trap almost always snaps. Comparing scaling-aware ratings for different departments, locations, or promotion types instead allows users to focus development efforts on those data subsets for which the gap to the ideal behavior is most pronounced. For example, when monitoring forecasting performance in time, usual metrics will change due to increasing or decreasing overall volume. Thresholds on scaling-unaware metrics (“if WMAPE increases above 70%, trigger an alarm!”) are prone to false positives and false negatives, which scaling-aware ratings remedy.
@@ 6.2 Future venues
Our proposed method – computing expectation values of metrics, and bucketing by similar predictions – can be applied to all probabilistic forecasts for which one can characterize the ideal case. Within the application in retail, several generalizations are thinkable: In a business context, it is advisable to judge the forecast by its actual economic impact (Syntetos et al., 2010), which can be rated in the context of what could be possibly achieved. A certain out-of-stock or waste loss, expressed in currency, can then be given a rating (“perfect”, “excellent”, etc.). The visibility of the financial impact of forecasting error, together with the knowledge about what could be achieved if the forecast were improved to a certain quality level, will help allocate resources efficiently. The question naturally emerges to which extent human overrides of forecasts (Khosrowabadi et al., 2022) intuitively incorporate precision scaling.
The exponent $\gamma=1.5$ in the parametrization in Eq. (11) works remarkably well and quite universally, for many different applications and industries (compare the M5-analysis to the Sainsbury’s analysis – models on both datasets follow the shape of the reference lines very well, across several orders of magnitude). Nevertheless, a better theoretical explanation or empirical corroboration of that value would be appreciated. Our hope is fuelled by the fact that scaling laws often reflect underlying system structures (West, 2018; Athanasopoulos and Kourentzes, 2022).
As a simple (one could even say: simplistic) first-order approximation to a benchmark, we used the Poisson distribution. For a second-order approximation, the distribution should be corrected for multi-buys (a single customer buys several items of the same product), possible finite capacity and other effects. The gap between empirical data and Poisson ideal in the bucketed representation of the Sainsbury’s dataset is, to a great extent, borne by multi-buy-effects.
For non-individualized retail, it is straightforward to identify the Poisson distribution as an ideal, maximally sharp distribution (Tichy, 2023c). For other processes, it might be much more involved to work out the sharpest possible forecast. Also, it is not guaranteed that the segregation by buckets of similar predicted means is sufficient: More in general, one will need to bucket in a way that “similar forecasts” (in terms of the distributions that they describe) are grouped and evaluated together. The size and number of the buckets should then ensure a good balance between precision (many small, homogenous buckets) and sufficient statistics per bucket (few large more inhomogeneous buckets). The general idea put forward in this article, segregating a large and heterogeneous group of forecasts into homogeneous buckets, is independent of the Poisson distribution.
On the conceptual side, it is worthwhile connecting the present work with the concepts of aleatoric and epistemic uncertainty (Hüllermeier and Waegeman, 2021): Any error metric value is the sum of an aleatoric, unavoidable error, and of an epistemic error, which may be decreased by incorporating further knowledge into the model. Using the Poisson distribution as baseline essentially states that the Poisson process describes the aleatoric part, while the possibly not perfectly known rate of that Poisson process is subject to epistemic uncertainty.
@@ Appendix A Heuristic algorithm for non-parametric Expectation Maximization
For the purpose of discussing our rating methodology on “perfect” Poissonian forecasts for a given dataset of sales values $s_{j}$, we construct predictions using the following heuristic, which is equivalent to Expectation Maximization: Given the set of observed sales $s_{j}$, we assume that these integer outcomes are the results of Poisson processes with rates $t_{j}$, i.e. that Eq. (4) holds. Given the known observations $s_{j}$, we need to evaluate $P(t|s)$, i.e. the conditional probability to have predicted $t$ given that we do observe $s$, to create plausible predictions $t_{j}$ (Tichy, 2023b). Remember that Eq. (4) states the conditional probability to observe $s$, given that $t$ is predicted. Bayes’ rule gives
$P(t|s)=\frac{P_{\text{Poi}}(s|t)P_{\text{rate}}(t)}{P_{\text{observation}}(s)},$ (12)
where $P_{\text{Poi}}(s|t)=\frac{e^{-t}{t^{s}}}{s!}$ is the Poisson distribution. The probability of a certain outcome $P_{\text{observation}}(s)$ can be easily retrieved from the dataset (it is the histogram of sales values). To evaluate $P(t|s)$, we need an approximation to $P_{\text{rate}}(t)$, the prior probability density function that describes the distribution of the rates.
We approximate this prior by starting with a first guess, an exponential distribution,
@@
where, intuitively speaking, we boost the probability density that contributes to those observations that are currently under-predicted. The updated probability distribution $\tilde{P}_{k+1}(t)$ is unnormalized, such that we perform the simple normalization
@@
To provide a set of plausible predictions $t_{j}$, we sample one prediction $t_{j}$ for each observation $s_{j}$ via $t_{j}\sim P(t|s_{j})$. Applying the iterative procedure on all data categories (product groups, locations…) separately, we obtain unbiased Poissonian forecasts that could have yielded the set of sales $s_{j}$. We have experienced that a dozen iterations is typically sufficient to reach a good artificial forecast, i.e. a forecast for which $P^{(k)}_{\text{observation}}(s)\approx P_{\text{observation}}(s)$ while, for each prediction $t$, the resulting sales $s$ are Poisson-distributed, as shown in Figure 1. In practice, we discretize the space of rates $t$ into a granular array of length $G=5^{\prime}000$, $t_{0}\dots t_{G}$, such that (16) is performed on all $P_{k}(t_{0})\dots P_{k}(t_{G})$ in a numerical fashion.
@@ A.1 Treatment of improbable observations
In the update rule (16), the cases $P^{(k)}_{\text{observation}}(s)=0$ and/or $P_{\text{observation}}(s)=0$ need special attention, especially in practice, when one recurs to numerical discretization of the space of rates $t$.

### v5

@@ 4.2 Demand Edge for Retail predictions at Sainsbury’s
The M5 dataset discussion in the previous section has the great advantage to be reproducible. On the other hand, one deals with a curated dataset that does not contain all data that are used in production (e.g., the price of the product). To complement our discussion and prove the applicability of our method in practice, we apply the scaling-aware rating to the predictions of the Demand Edge for Retail forecasts produced by Blue Yonder for the UK retailer Sainsbury’s. Demand Edge for Retail is based on a causal machine-learning model that computes the contribution of features such as price, weather, day-of-the-week, holiday, promotion and many others to the expected demand of a product (Wick et al., 2019, 2021). It produces a forecast that consists of a prediction for the expected mean as well as for the variance, technical details can be found in (Wick, 2023a, b, c). To simplify the discussion and make it comparable to the M5 dataset discussion, we only use the predicted mean, and interpret it as predicted mean of the Poisson distribution.
@@
We analyse the next-day-forecasts (horizon 1) for Monday, October 2nd to Saturday, October 28th, 2023, excluding Sundays. We exclude items that are not sold in units of pieces but volume (liters) or mass (KGs), predictions that are smaller than 0.1 / product-location-day, and unsold product-day-location combinations for which the prediction was 20 or larger, for which we can safely assume an out-of-stock-situation. After data cleaning, we remain with more than 250 million product-location-day combinations, corresponding to more than 500 million sold items. Global model over-prediction is less than 1%.
@@
Figure 8 shows the bucketed scatter plots for bias and noise for the Sainsbury’s dataset. It confirms the calibration of the forecast (outside the ultra-slow-movers), and the only slightly super-Poissonian noise over three orders of magnitude. Our approach confirms that, on a global level, the prediction is comparable to what could possibly be expected from first principles (Tichy, 2023c), as discussed in Section 2.1. This is also reflected by the excellent values $S_{\text{overall}}(\text{NMRPS})=84.7\%$ and $S_{\text{overall}}(\text{Bias})=78.7\%$.
@@
A more in-depth investigation confirms that the reasons for non-Poissonian behavior described in Section 2.2 above are indeed relevant in this scenario (there are, for example, substantial “multi-buy” events when customers buy several pieces of one item). A thorough quantitative understanding of the mild departure from the Poisson ideal case based on those effects would be highly appreciated, but remains beyond the scope of the current work.
@@
One can segregate data by different dimensions to ask questions as the following: Are certain product groups/stores/weekdays predicted better than others? All of these comparisons risk to induce the naïve scaling-trap, which is resolved by our approach. As an example, we differentiate by the highest-level product group and focus on the three largest assortments (Fresh, Ambient, Non foods). The resulting metrics are shown in Table 5.
@@
Together with Figure 9, we see that the overall numeric metrics (e.g., WMAPE, NMRPS) are misleading if used directly, since Fresh tends to sell at about twice the average velocity – one more example of the naïve scaling trap. Once we apply our bucketed evaluation summarized by Table 6, we see that all three groups maintain comparable, near-Poisson performance, earning “Good” or “Excellent” ratings. We use the same parameters for ratings as for the M5 competition to make the numbers comparable; Blue Yonder uses internally different sets of industry-specific parameters.
@@
Table 6: Scaling-aware ratings for the three main product groups at Sainsbury’s, for Blue Yonder’s Luminate Demand Edge forecasting solution, summarizing the prediction-bucketed representation of Figure 9.
@@
Figure 9: Prediction-bucketed representation of noise (NMRPS), for Demand Edge for Retail predictions by Blue Yonder for UK retailer Sainsbury’s, differentiated by the highest-level product group. Circle sizes reflect the square-root of total number of sales contained in the respective buckets.
@@
These examples show that applying scaling-aware rating on different data-subsets such as departments, locations, days-of-the week helps focus development efforts on those specific forecasts that exhibit the largest gap to the Poisson ideal case. These are often not those that exhibit extremal values of scaling-unaware metrics.
@@ 5 Conclusions and outlook
Applied forecast judgement methodology should reconcile prevailing business practice (Carvallo, 2021; Bower, 2023) with the state of the art in probabilistic forecasting (Czado et al., 2009; Gneiting, 2011; Kolassa, 2016; Wei and Held, 2014; Gneiting and Katzfuss, 2014). Scaling-aware forecast rating provides a framework to tackle the problems of benchmarking and precision scaling in a systematic way. By adapting the parametrization introduced in Section 3.2, one can apply scaling-aware forecast rating to different forecasting problems, e.g. to industries in which bad data quality is common or certain influencing factors are known to be uncontrollable.
@@ 5.1 Applicability
Given a forecast, our method permits users to judge the status quo and whether desired improvements in accuracy are realistic at all: If forecast performance is close to the ideal Poisson prediction (or another distribution known to be the sharpest possible for the task at hand), the ask for a substantial improvement needs to be refused. If a gap to the Poissonian case emerges, the reason for this gap needs to be understood. Forecast creators should then elaborate whether it is a realistic expectation to push the forecast closer to the Poisson ideal.
When several models compete, using scaling-aware methods makes their comparison more transparent and insightful: A given improvement of a metric can be understood holistically by analysing the respective bucket scatter plot as in Figures 6, 7, 8, 9. Two competing models can be placed onto a meaningful scale from “perfect” to “unacceptable”, providing context and orientation whether a certain improvement is substantial and solid, or little more than a mere coincidence.
The greatest added value of our method lies in dataset comparison: A comparison of the values of scaling-unaware metrics such as MAE, WMAPE or (N)MRPS for two sub-datasets of the same model is useless, since the naïve scaling trap almost always snaps. Comparing scaling-aware ratings for different departments or locations instead allows users to focus development efforts on those data subsets for which the gap to the ideal behavior is most pronounced. For example, when monitoring forecasting performance in time, usual metrics will change due to increasing or decreasing overall volume. Thresholds on scaling-unaware metrics (“if WMAPE increases above 70%, trigger an alarm!”) are prone to false positives and false negatives, which scaling-aware ratings remedy.
@@ 5.2 Future venues
Our proposed method – computing expectation values of metrics, and bucketing by similar predictions – can be applied to all probabilistic forecasts for which one can characterize the ideal case. Within the application in retail, several generalizations are thinkable: In a business context, it is advisable to judge the forecast by its actual economic impact (Syntetos et al., 2010), which can be rated in the context of what could be possibly achieved. A certain out-of-stock or waste loss, expressed in currency, can then be given a rating (“perfect”, “excellent”, etc.). The question naturally emerges to which extent human overrides of forecasts (Khosrowabadi et al., 2022) intuitively incorporate precision scaling.
The exponent $\gamma=1.5$ in the parametrization in Eq. (10) works remarkably well and quite universally, for many different applications and industries (compare the M5-analysis to the Sainsbury’s analysis – models on both datasets follow the shape of the reference lines across several orders of magnitude). Nevertheless, a better theoretical explanation of that value would be appreciated. Our hope is fuelled by the fact that scaling laws often reflect underlying system structures (West, 2018; Athanasopoulos and Kourentzes, 2022).
In the case of finite capacity, which can result in stock-outs, one can bucket the data both by the predicted uncensored selling rate and by the predicted probability to go into stock-out Tichy (2024c). This then allows again to compare the expected value of the metric to the achieved one.
@@ Acknowledgements
Data Availability
The data of the M5 competition that support the findings of this study are openly available at kaggle (2022a). Blue Yonder and Sainsbury’s did not give written consent for their data to be shared publicly.
@@ Appendix A Heuristic algorithm for non-parametric Expectation Maximization
To discuss our rating methodology on “perfect” Poissonian forecasts for a given dataset of sales values $s_{j}$, we construct baseline predictions using the following heuristic, which is equivalent to Expectation Maximization: Given the observed sales $s_{j}$, we assume that these integer outcomes are the results of Poisson processes with rates $t_{j}$, i.e., that Eq. (3) of the main text holds. Given the known observations $s_{j}$, we need to evaluate $P(t|s)$, i.e., the conditional probability to have predicted $t$ given that we do observe $s$, to create plausible predictions $t_{j}$. Remember that Eq. (3) of the main text states the conditional probability to observe $s$, given $t$. Bayes’ rule gives
$P(t|s)=\frac{P_{\text{Poi}}(s|t)P_{\text{rate}}(t)}{P_{\text{observation}}(s)}.$ (11)
The probability of a certain outcome $P_{\text{observation}}(s)$ is the empirical frequency of that value $s$ in the dataset. To evaluate $P(t|s)$, we need an approximation to $P_{\text{rate}}(t)$, the prior probability density distribution of the rates.
We approximate this prior by starting with a an exponential distribution,
@@
where, intuitively speaking, we boost the probability density that contributes to those observations that are currently under-predicted. The updated probability distribution $\tilde{P}_{k+1}(t)$ is unnormalized, such that we apply
@@
To provide a set of plausible predictions $t_{j}$, we sample one prediction $t_{j}$ for each observation $s_{j}$ via $t_{j}\sim P(t|s_{j})$. Applying the iterative procedure on all data categories (product groups, locations…) separately, we obtain unbiased Poissonian forecasts that could have yielded the set of sales $s_{j}$. We have experienced that a dozen iterations is typically sufficient to reach a good artificial forecast, i.e., a forecast for which $P^{(k)}_{\text{observation}}(s)\approx P_{\text{observation}}(s)$ while, for each prediction $t$, the resulting sales $s$ are Poisson-distributed (see Figure 10). In practice, we discretize the space of rates $t$ into a granular array of length $G=5^{\prime}000$, $t_{0}\dots t_{G}$, such that (15) is performed on all $P_{k}(t_{0})\dots P_{k}(t_{G})$ in a numerical fashion.
@@ A.1 Treatment of improbable observations
In the update rule (15), the cases $P^{(k)}_{\text{observation}}(s)=0$ and/or $P_{\text{observation}}(s)=0$ need special attention.
@@
A.2 Application to M5 dataset
The calibration diagrams of Figure 10 confirm that the baseline $\vec{r}^{\text{baseline}}$ that we constructed by following the EM procedure is an excellent baseline model, since the unconditioned and the conditioned distributions of actually observed sales $s$ match the Poisson prediction well.
Figure 10: Calibration diagrams for ideal baseline model, based on the heuristic of A: Distribution of observed sales frequency (orange) vs. predicted probability (blue, this color refers to the ideal Poisson case throughout this article). Upper left: Overall marginal calibration diagram (for visual convenience, we mapped the observation 0 to 0.5 to fit into the log-log-scale). Upper right: Cut on $1.8<r<2.2$. Lower left: Cut on $4.8<r<5.2$. Lower right: Cut on $9.8<r<10.2$.
Appendix B Summarizing buckets to overall ratings
We elaborate on the summary of bucket-wise to overall ratings that we sketched in Section 3.4 of the main text. We assume to hold a set of prediction-stratified buckets indexed by $R$, the rounded logarithm of the rate defined. Each bucket contains $n_{R}$ prediction/observation pairs, denoted by $\vec{r}^{(R)},\vec{s}^{(R)}$. For each bucket, we compute the total and the average prediction
$r_{\text{total}}^{(R)}$ $=$ $\sum_{j,R_{j}=R}r_{j},$ (17)
$r_{\text{mean}}^{(R)}$ $=$ $r_{\text{total}}^{(R)}/n_{R},$ (18)
the total and the average observation
$s_{\text{total}}^{(R)}$ $=$ $\sum_{j,R_{j}=R}s_{j},$ (19)
$s_{\text{mean}}^{(R)}$ $=$ $s_{\text{total}}^{(R)}/n_{R}$ (20)
and the achieved metrics $M^{(R)}_{\text{actual}}=M(\vec{r}^{(R)},\vec{s}^{(R)})$. By plotting the achieved metrics against the predicted rates in the context of the metrics reference values (i.e., using Figure 5 of the main text as a background context), one can visually extract useful insights on model quality and possible improvements, as exemplified in Section 4 of the main text.
In certain situations, it might be desirable to condense and summarize the information contained in a set of buckets by a single number. There are two possibilities to perform this summary: aggregate the ratings of the metric per bucket to yield an overall score, or judge the overall aggregated metric, e.g. the overall achieved WMAPE.
B.1 Aggregate bucket ratings
Given an evaluation metric $M$, we compute the ratings of the individual buckets and average them to an overall rating: For each bucket $R$, the achieved metric value $M^{(R)}_{\text{actual}}$ is compared to the reference values of the metric for the different ratings, i.e., to
$M^{(R)}_{\text{perfect}},~M^{(R)}_{\text{excellent}},\dots,M^{(R)}_{\text{unacceptable}}$. (21)
Let the metric $M$ be a loss function, such that larger values are worse, and $M^{(R)}_{\text{actual}}>M^{(R)}_{\text{unacceptable}}$ reveals some dramatic problem of the forecast in the respective bucket.
To define numerical scores that allow an average of the ratings resulting in some overall score $S_{\text{overall}}$, we map quality categories (“perfect”, “excellent”, …,“unacceptable”) to percentage ratings $S_{\text{quality}}$, using a scale between 100% (best) to 0% (worst):
Quality Perfect Excellent Good OK Fair Insufficient Unacceptable
$S_{\text{quality}}$ $91.67\% >$$75\% >$$58.33\% >$$41.67\% >$$25\% >$$\% <0.083<0.083\%$
For each bucket, the achieved metric $M^{(R)}_{\text{actual}}$ implies some $M^{(R)}_{\text{upper}}<M^{(R)}_{\text{actual}}<M^{(R)}_{\text{lower}}$ defining the two neighboring qualities, with percentage ratings $S_{\text{upper}}>S_{\text{lower}}$. For the bias, we map negative bias values $r_{\text{total}}^{(R)}/s_{\text{total}}^{(R)}=b<1$ to $1/b>1$ to achieve the monotonicity of our rating (for example, we rate a bias of 1.1 as being equivalent to a bias of $1/1.1\approx 0.91$).
We set the percentage rating of the bucket to the linear interpolation of the ratings achieved by the neighboring buckets:
$S^{(R)}=S_{\text{lower}}+(S_{\text{upper}}-S_{\text{lower}})\frac{M^{(R)}_{\text{actual}}-M^{(R)}_{\text{upper}}}{M^{(R)}_{\text{lower}}-M^{(R)}_{\text{upper}}}$, (22)
where this pragmatic choice might be replaced by splines or other interpolation methods. Eq. (22) assigns a numerical score to each bucket. A score of 100% is achieved when the metric computed in the bucket matches exactly the value expected under the Poisson distribution. A score of 0.083% is awarded when the metric matches the value $M^{(R)}_{\text{unacceptable}}$ that is achieved by a distribution affected by “unacceptable” noise (according to the rating). The worst score, 0%, is reached when twice the “unacceptable” value is attained for loss functions. The numerical score thus quantifies “how far from the ideal” the bucket lies in the prediction-metric-area, and how that shall be interpreted operationally.
The overall rating $S_{\text{overall}}(\vec{r},\vec{s})$ is defined by the weighted mean of the $S^{(R)}$ that are awarded to the buckets:
$S_{\text{overall}}(\vec{r},\vec{s})=\frac{\sum_{R}S^{(R)}\cdot\text{max}(s^{(R)}_{\text{total}},r^{(R)}_{\text{total}})}{\sum_{R}\text{max}(s^{(R)}_{\text{total}},r^{(R)}_{\text{total}})}$ (23)
where each bucket is weighted by either its total prediction or its total observation, depending on which is larger. We do not merely weight by total bucket sales or by total bucket predictions, but by bucket-wise maximum to avoid neglecting buckets with extreme over- or under-forecasting. The resulting $S_{\text{overall}}$ is then the scaling-aware rating of the overall set of predictions and observations, which can corroborate operational decisions.
B.2 Rate the overall aggregated metric
The overall metric $M(\vec{r},\vec{s})$ is embedded into a context by computing reference values $M_{\text{perfect}}(\vec{r})$, $M_{\text{excellent}}(\vec{r})$, $M_{\text{unacceptable}}(\vec{r})$ as follows: For a given set of buckets, we compute the overall metric $M_{\text{quality}}(\vec{r})$ which would have been achieved if all buckets $R$ had achieved the $M^{(R)}_{\text{quality}}$ from Eq. (21). This allows the user to corroborate statements such as “an overall WMAPE of 44% was achieved, which is between good (38%) and OK (46%)”.
The rating of the overall aggregated metric and the weighted mean rating of the buckets will often yield similar conclusions. They are, however, answering a different question and can, in certain cases, deviate substantially: Consider a forecast that is unbiased overall,
$\sum_{R}r^{(R)}_{\text{total}}\approx\sum_{R}s^{(R)}_{\text{total}}$. (24)
The overall aggregated bias is, thus, excellent. However, it might still be the case that slow-movers are heavily underforecasted while fast-movers are overforecasted (or vice versa), such that
$\forall R:r^{(R)}_{\text{total}}\gg s^{(R)}_{\text{total}}\text{ or }r^{(R)}_{\text{total}}\ll s^{(R)}_{\text{total}}$ (25)
The scaling-aware rating then yields a critical value (such as “fair”, “insufficient”…) because each bucket is biased, while the overall aggregated bias seems to be excellent.